A person drops a cylindrical steel bar ( Y = 6.00 × 10 10 Pa ) from a height of 1.10 m (distance between the floor and the bottom of the vertically oriented bar). The bar, of length L = 0.920 m , radius R = 0.00500 m , and mass m = 1.300 kg , hits the floor and bounces up, maintaining its vertical orientation. Assuming the collision with the floor is elastic, and that no rotation occurs, what is the maximum compression of the bar?
Answer #1

Calculate the velocity of the bar when it hits the ground using equations of motion:

v^2 = vo^2 + 2 a d

v^2 = 2 a d

In elastic collision the energy is conserved. the kinetic energy just before it hits the ground should be equal to the stored potential energy when it is fully compressed.

KE = PE

Calculate kinetic energy,

KE = 1/2 m v^2

KE = 1/2 m *2* a *d

KE = m * a * d

Spring potential energy is

PE = 1/2 * k* x^2.

Write the spring force,

F = k*x.

Using the definition of Young's modulus and some algebra you can see that k = F/x = YA/L.

The potential energy becomes,

PE = 1/2 k x^2

= 1/2 Y A x^2 / L

= 1/2 Y A x^2 / L

PE = 1/2 * Y * π * R^2 * x^2 / L

According to the conservation of energy, potential energy equal to kinetic energy and solve for x:

PE = KE

1/2 Y* π* R^2 x^2 / L = m a d

x = sqrt{2 m *a* d* L / [Y* π* R^2]}

= sqrt{2 (1.3 kg) (- 9.81 m/s^2) (-1.10 m) (0.92 m) / [(6 * 10^10 Pa) pi (0.005 m)^2]}

= 2.34 * 10^-3 m

x = 2.34 mm

Thus, the maximum compression in the bar is 2.34 mm.